# Dark matter cooling during early matter-domination boosts sub-earth halos

### v1

@@ 3.1 Early matter domination: temperature dependent entropy injection
We consider a background cosmological model where the oscillating scalar field $\phi$, with equation of state $\omega_{\phi}=0$, dominates the universe prior to the BBN and dissipates its energy into the radiation at a rate
@@
where $\Gamma_{0}$ is a dimensionless constant. The time evolution of the energy densities of the meta-stable scalar field ($\rho_{\phi}$) and its decay products ($\rho_{\gamma}$), and the number density of the DM ($n_{\chi}$) are governed by the Boltzmann equations given below
@@
We consider that the DM remains non-relativistic at all times, thus we take $m_{\chi}\geq T_{\rm max}$. Despite initially having $n_{\chi}(0)=0$, the DM rapidly reaches equilibrium with the plasma owing to its substantial interactions with the bath particles, and eventually undergoes chemical freeze-out around $T_{\rm fo}\sim m_{\chi}/20$. For illustration, we sketch the evolution of the background energy densities $\rho_{\phi}$ and $\rho_{\gamma}$ scaled by the total energy density in Fig. 2, using the full numerical solutions of the set of Boltzman Eqs. (3.2), (3.3), and (3.4). In Fig. 2, we also show the evolution of the DM number density $n_{\chi}$ and compare it with the equilibrium number density $n_{\chi,{\rm eq}}$ for the Model I described in Sec. 4.1.
@@ 3.2 Evolution of matter perturbations
We are mainly interested in the modes (with wave number $k$) which are super-Hubble, i.e. $k\ll aH$ at $a_{\rm max}$, but subsequently enters the horizon during the EMDE epoch. Since the background energy density of the universe is dominated by $\phi$ at the onset of EMDE, we impose adiabatic initial conditions at $a_{\rm max}$ for the multi-fluid perturbations such that
@@
In Fig. 3, we present the numerical solutions of the perturbation Eqs. (A.1) to (A.10), in Fourier space as a function of $a/a_{\rm RH}$ using a benchmark parameter set. In the left panel, the evolution of $\Phi$, $\delta_{\gamma}$, and $\theta_{\gamma}$ are shown for a mode $k=30\;k_{\rm RH}$, where $k_{\rm RH}=a_{\rm RH}H(a_{\rm RH})$ represents the wave number of the mode that enters the horizon at $T_{\rm RH}$. The $k=30\;k_{\rm RH}$ mode enters the horizon at $a\simeq 1.9\times 10^{-3}a_{\rm RH}$ while the reheating concludes at $a_{\rm RH}/a_{\rm in}=1.7\times 10^{7}$. We observe that $\Phi$ remains nearly constant during the EMDE, while $\delta_{\gamma}$ receives an initial kick during the horizon crossing and starts to grow due to the source term arising from the dissipation of $\phi$, see Eq. (A.3). However, this growth in $\delta_{\gamma}$ does not sustain till the end of EMDE due to a back-reaction from the growing spatial dispersion (increasing $\theta_{\gamma}$) of the radiation fluid. All in all, $\delta_{\gamma}$ reaches a peak and then decayed before the reheating ends, and oscillates with a suppressed amplitude once the radiation domination begins.
@@
The density perturbation of the DM ($\delta_{\chi}$) follows the $\delta_{\chi,{\rm eq}}$ as long as the DM remains in chemical equilibrium. However, after the freeze-out ($a>a_{\rm fo}$), $\delta_{\chi}$ decreases to $-2\Phi_{0}$ and remains constant until the corresponding mode enters the horizon. This is consistent with the adiabatic initial condition in Eq. (3.10), since $\rho_{\phi,\chi}\sim a^{-3}$ after freeze-out. After the horizon entry of the mode ($a>a_{\rm hor}$), $\delta_{\chi}$ grows linearly until reheating, followed by a logarithmic scaling during radiation domination, and resumes linear growth after matter-radiation equality for $a>a_{\rm eq}$. The leading scaling behaviour of $\delta_{\chi}$ with the scale factor at different epochs are summarized below
@@ 3.3 Matter power spectrum
Formation of the small scale structures due to the linear growth of matter perturbations for the modes that become sub-Hubble during the EMDE start well after the universe went through the matter-radiation equality. The effect of the EMDE enters into the matter power spectrum through a mode-dependent rescaling of the transfer function $T(k)$ as $T(k)\to R(k)T(k)$, where $R(k)$ accounts for the boost in the matter power spectrum due to the linear growth of perturbations during the EMDE.
@@
Here $\Omega_{\chi}$ and $\Omega_{b}$ denotes the relic density of the DM and the baryons, respectively. The scale factor at the horizon entry $a_{\rm hor}(k)$ for modes with different scales are given by
@@
Figure 4: The matter power spectrum $P_{\chi}(k)$ for different values of $\lambda_{\rm fsh}$, evaluated at $z=50$, in case of an EMDE with a temperature dependent entropy injection, where $T_{\rm RH}=10$ MeV, and $k_{\rm RH}=1.7\times 10^{7}\;k_{\rm eq}$. We use the same set of benchmark points as mentioned in Fig. 3. The green dot-dashed line represents $P_{\chi}(k)$ for purely radiation dominated universe with $\lambda_{\rm fsh}=0$. The coloured vertical lines correspond to the modes for which $k\sim\lambda_{\rm fsh}^{-1}$, and have the values $k_{\rm fsh}=3.8\times 10^{8}\;k_{\rm eq}$ and $5.3\times 10^{9}\;k_{\rm eq}$, for Model I and Model II, respectively. The matter power spectrum $P_{\chi}(k)$ shows different scaling behavior on two sides of $k_{\rm RH}$ following Eq. (3.19).
@@
The solid red curve in Fig. 4 presents the matter power spectrum in the presence of EMDE with energy dependent dissipation. The leading momentum dependence of $P_{\chi}(k)\propto|\delta_{\chi}|^{2}$ for the modes entering the horizon before and after reheating is given by
@@ 3.4 Formation of sub-earth halos
The enhancement of power spectrum for the modes $k_{\rm RH}<k<k_{\rm fsh}$ dictate the range of mass scales for the DM halos having enhanced population compared to the standard radiation dominated universe. Two characteristic mass scales, setting the limits on the masses contained in the halos with enhanced populations are
@@
Left panel of the Fig. 5 presents the variation of $\sigma(M)$ with the halo mass $M$ normalized to the mass of the earth ($M_{\oplus}$) for different values of the free-streaming horizon, at a redshift $z=10$, and $T_{\rm RH}=10$ MeV. Clearly, for $M>M_{\rm RH}$, $\sigma(M)$ follows the rms density that would be found in a standard radiation dominated universe with $\lambda_{\rm fsh}=0$. However, for $M_{\rm fsh}<M<M_{\rm RH}$, a power law growth $\sigma\left(M\right)\propto M^{-\left(n_{s}+3\right)/6}$ is observed, which becomes constant at scales below $M_{\rm fsh}$.
Figure 5: $\sigma(M)$ (left) and differential mass fraction (right) at $z=10$ are shown as function of dark matter halo mass in units of earth mass($M_{\oplus}$), for different values of $\lambda_{\rm fsh}$. We set $T_{\rm RH}=10$ MeV which corresponds to $M_{\rm RH}=10.5\,M_{\oplus}$. The vertical dotted contours denote the value of $M_{\rm fsh}$ associated with the respective free-streaming horizons. The green dot-dashed curves represent the scenario with pure radiation domination and $\lambda_{\rm fsh}=0$.
@@
where $\delta_{c}\simeq 1.686$ is the critical value of linear overdensity for $z\geq 2$. In the right panel of Fig. 5 we show that the increase in the differential fraction significantly depends on the free-streaming horizon of the DM. In fact, we observe that if $\lambda_{\rm fsh}$ is below a threshold value $\lambda_{\rm fsh}^{c}$, the structures forming at scales smaller than $M_{\rm RH}$ in the presence of EMDE are guaranteed to show a boost in the differential fraction compared to the radiation dominated universe. The threshold $\lambda_{\rm fsh}^{c}\simeq 1.6$ pc primarily depends on $T_{\rm RH}$. Since we determine the threshold $\lambda_{\rm fsh}^{c}$ by comparing with a scenario where $\lambda_{\rm fsh}=0$ in the purely radiation dominated universe, any $\lambda_{\rm fsh}<\lambda_{\rm fsh}^{c}$ in the case of EMDE ensures a boost in the structure formation.
Enhanced small-scale structures could amplify DM annihilation signals within sub-earth halos, providing a probe for a variety of DM models as well as non-standard cosmological epochs like EMDE. The dynamics of DM decoupling during EMDE determine the free-streaming horizon, which, when compared to the threshold horizon $\lambda_{\rm fsh}^{c}$, can help identify viable range of microscopic model parameters compatible with enhanced structures, we defer further discussion to Sec. 4.
@@ 4.1 Model I: s-wave scattered scalar dark matter
The first model we consider has a scalar dark matter ($\phi_{\chi}$), interacting with scalar bath particles ($\phi_{\gamma}$) via a quartic vertex $(\lambda/4)\phi_{\chi}^{2}\phi_{\gamma}^{2}$. In general, $\phi_{\gamma}$ may represent either the Standard Model Higgs boson, or any BSM scalar in thermal bath. However, in the following, we assume that the mass of $\phi_{\gamma}$ is much smaller than all the relevant scales. This leads to an $s$-wave scattering $\phi_{\chi}\phi_{\gamma}\to\phi_{\chi}\phi_{\gamma}$, with momentum transfer rate
@@
where $m_{\chi}\in(20-1000)$ GeV is the typical mass for an WIMP DM.
@@
Figure 7: The allowed parameter space satisfying the present day relic density of the DM is shown with distinct values of $T_{\rm RH}$ presented by coloured bands. The DM shows partial kinetic decoupling during the EMDE in the dark shaded region, while in the light shaded region DM kinetically decouples after reheating. In the lime coloured region below the colour bands, DM produces non-thermally from the bath, while the light cyan region, where $T_{\rm RH}<10$ MeV, is excluded to avoid bounds from the BBN. The isocontours of the free-streaming horizon are shown by the white curves. Amplification of small scale structures compared to the standard radiation dominated cosmology occurs to the right of the cyan dot-dashed contour, which represent the cut-off value of the free-streaming horizon $\lambda^{c}_{\rm fsh}$.
@@
  • The DM was in thermal contact at the onset of the EMDE, and freezes-out prior to the end of reheating, the observed present day relic density of the DM $\Omega^{0}_{\rm DM}h^{2}=0.12$ [2] is reproduced in the entire coloured region.
  • In the dark coloured area, partial kinetic decoupling of the DM occurs, however, the sub-earth halo population amplifies below the cyan dot-dashed contour.
@@
A few comments regarding the impact of the temperature dependent entropy injection during the EMDE is in order. First of all, the annihilation cross-section required to satisfy the relic density in this cosmological background is smaller than that required in a fully radiation dominated universe, due to the dilution of the relic by the entropy injection. In fact, in the parameter space of Fig. 7 that show freeze-out production of the DM, $\langle\sigma v\rangle_{\rm ann}$ ranges from $10^{-11}-10^{-15}$ GeV-2 depending on $m_{\chi}$, which is at least two orders of magnitude smaller than what would be required in the absence of EMDE. This reduction in the annihilation cross-section helps evade the strong bounds coming from the direct detection experiments. On the other hand, we emphasize that the required cross-section in this scenario is in principle larger than that in an EMDE with constant rate of entropy injection ($\Gamma_{\phi}\sim{\rm const.}$). To give an idea, we estimate the ratio of $\langle\sigma v\rangle_{\rm ann}$, required to produce the same comoving number density of the DM after reheating in EMDEs with constant and temperature dependent entropy injection rates as follows
@@
Here $x^{T}_{\rm fo}$ and $x^{c}_{\rm fo}$ denote the ratio $m_{\chi}/T_{\rm fo}$ in the two scenarios with $\Gamma_{\phi}\sim T$ and $\Gamma_{\phi}\sim{\rm const.}$, respectively. Assuming that the freeze-out occurs around $x_{\rm fo}\sim 20$ in both scenarios and neglecting the $T/m_{\chi}$ dependence in $\langle\sigma v\rangle_{\rm ann}$, we find that the annihilation cross-section required for the same amount of relic is approximately 50 times larger in the $\Gamma_{\phi}\sim T$ case compared to the $\Gamma_{\phi}\sim{\rm const.}$ case. However, we also point out that the present day relic density of the DM, $\Omega^{0}_{\rm DM}h^{2}=0.12$ [2], cannot be satisfied by the quartic interaction during an EMDE with $\Gamma_{\phi}\sim{\rm const.}$, if $T_{\rm RH}\ll 10$ GeV [40]. Therefore, the temperature dependent entropy injection accommodates a novel region of the parameter space where the relic density can be satisfied even with low reheating temperature $T_{\rm RH}\sim 100$ MeV, and reduced annihilation cross-section compared to the radiation dominated universe.
The above discussion indicates the crucial impact of the background cosmological epoch on the production of the DM. Moreover, in a large range of $m_{\chi}$, the DM with $s$-wave scattering shows partial kinetic decoupling during the EMDE. This is also unique for $\Gamma_{\phi}\sim T$ scenario, as $s$-wave partial amplitudes are not sufficient to kinetically decouple the DM in $\Gamma_{\phi}\sim{\rm const.}$ scenario, see Table 1. Thus, a broad region of the parameter space exhibits the correct freeze-out relic density with a low reheating temperature, leading to a reduced DM free-streaming horizon below the cut-off $\lambda^{c}_{\rm fsh}$. The resulting enhancement in small-scale structure formation at sub-earth mass scales offers new avenues to probe this scenario through DM annihilation signatures, sub-halo DM interactions with cosmic rays, and sub-halo collisions with neutron stars.

### v2

@@ 3.1 Early matter domination: temperature dependent entropy injection
We consider a background cosmological model where the oscillating scalar field $\phi$, with the equation of state $\omega_{\phi}=0$, dominates the universe prior to the BBN and dissipates its energy into the radiation at a rate
@@
where $\Gamma_{0}$ is a dimensionless constant. The time evolution of the energy densities of the meta-stable scalar field ($\rho_{\phi}$) and its decay products ($\rho_{\gamma}$) and the number density of the DM ($n_{\chi}$) are governed by the Boltzmann equations given below
@@
We consider that the DM remains non-relativistic at all times, thus we take $m_{\chi}\geq T_{\rm max}$. Despite initially having $n_{\chi}(0)=0$, the DM rapidly reaches equilibrium with the plasma owing to its substantial interactions with the bath particles and eventually undergoes chemical freeze-out around $T_{\rm fo}\sim m_{\chi}/20$. For illustration, we sketch the evolution of the background energy densities $\rho_{\phi}$ and $\rho_{\gamma}$ scaled by the total energy density in Fig. 2, using the full numerical solutions of the set of Boltzman Eqs. (3.2), (3.3), and (3.4). In Fig. 2, we also show the evolution of the DM number density $n_{\chi}$ and compare it with the equilibrium number density $n_{\chi,{\rm eq}}$ for the Model I described in Sec. 4.1.
@@ 3.2 Evolution of matter perturbations
We are mainly interested in the modes (with wave number $k$) which are super-Hubble, i.e. $k\ll aH$ at $a_{\rm max}$, but subsequently enter the horizon during the EMDE epoch. Since the background energy density of the universe is dominated by $\phi$ at the onset of EMDE, we impose adiabatic initial conditions at $a_{\rm max}$ for the multi-fluid perturbations such that
@@
In Fig. 3, we present the numerical solutions of the perturbation Eqs. (A.1) to (A.10), in Fourier space as a function of $a/a_{\rm RH}$ using a benchmark parameter set. In the left panel, the evolution of $\Phi$, $\delta_{\gamma}$, and $\theta_{\gamma}$ are shown for a mode $k=30\;k_{\rm RH}$, where $k_{\rm RH}=a_{\rm RH}H(a_{\rm RH})$ represents the wave number of the mode that enters the horizon at $T_{\rm RH}$. The $k=30\;k_{\rm RH}$ mode enters the horizon at $a\simeq 1.9\times 10^{-3}a_{\rm RH}$ while the reheating concludes at $a_{\rm RH}/a_{\rm in}=1.7\times 10^{7}$. We observe that $\Phi$ remains nearly constant during the EMDE, while $\delta_{\gamma}$ receives an initial kick during the horizon crossing and starts to grow due to the source term arising from the dissipation of $\phi$, see Eq. (A.3). However, this growth in $\delta_{\gamma}$ does not sustain till the end of EMDE due to a back-reaction from the growing spatial dispersion (increasing $\theta_{\gamma}$) of the radiation fluid. All in all, $\delta_{\gamma}$ reaches a peak and then decays before the reheating ends, and oscillates with a suppressed amplitude once the radiation domination begins.
@@
The density perturbation of the DM ($\delta_{\chi}$) follows the $\delta_{\chi,{\rm eq}}$ as long as the DM remains in chemical equilibrium. However, after the freeze-out ($a>a_{\rm fo}$), $\delta_{\chi}$ decreases to $-2\Phi_{0}$ and remains constant until the corresponding mode enters the horizon. This is consistent with the adiabatic initial condition in Eq. (3.10), since $\rho_{\phi,\chi}\sim a^{-3}$ after freeze-out. After the horizon entry of the mode ($a>a_{\rm hor}$), $\delta_{\chi}$ grows linearly until reheating, followed by a logarithmic scaling during radiation domination, and resumes linear growth after matter-radiation equality for $a>a_{\rm eq}$. The leading scaling behavior of $\delta_{\chi}$ with the scale factor at different epochs are summarized below
@@ 3.3 Matter power spectrum
Formation of the small-scale structures due to the linear growth of matter perturbations for the modes that become sub-Hubble during the EMDE start well after the universe went through the matter-radiation equality. The effect of the EMDE enters into the matter power spectrum through a mode-dependent rescaling of the transfer function $T(k)$ as $T(k)\to R(k)T(k)$, where $R(k)$ accounts for the boost in the matter power spectrum due to the linear growth of perturbations during the EMDE.
@@
Here $\Omega_{\chi}$ and $\Omega_{b}$ denote the relic density of the DM and the baryons, respectively. The scale factor at the horizon entry $a_{\rm hor}(k)$ for modes with different scales are given by
@@
Figure 4: The matter power spectrum $P_{\chi}(k)$ for different values of $\lambda_{\rm fsh}$, evaluated at $z=50$, in case of an EMDE with a temperature-dependent entropy injection, where $T_{\rm RH}=10$ MeV, and $k_{\rm RH}=1.7\times 10^{7}\;k_{\rm eq}$. We use the same set of benchmark points as mentioned in Fig. 3. The green dot-dashed line represents $P_{\chi}(k)$ for a purely radiation-dominated universe with $\lambda_{\rm fsh}=0$. The coloured vertical lines correspond to the modes for which $k\sim\lambda_{\rm fsh}^{-1}$, and have the values $k_{\rm fsh}=3.8\times 10^{8}\;k_{\rm eq}$ and $5.3\times 10^{9}\;k_{\rm eq}$, for Model I and Model II, respectively. The matter power spectrum $P_{\chi}(k)$ shows different scaling behavior on two sides of $k_{\rm RH}$ following Eq. (3.19).
@@
The solid red curve in Fig. 4 presents the matter power spectrum in the presence of EMDE with energy-dependent dissipation. The leading momentum dependence of $P_{\chi}(k)\propto|\delta_{\chi}|^{2}$ for the modes entering the horizon before and after reheating is given by
@@ 3.4 Formation of sub-earth halos
The enhancement of power spectrum for the modes $k_{\rm RH}<k<k_{\rm fsh}$ dictates the range of mass scales for the DM halos having an enhanced population compared to the standard radiation-dominated universe. Two characteristic mass scales, setting the limits on the masses contained in the halos with enhanced populations are
@@
The left panel of Fig. 5 presents the variation of $\sigma(M)$ with the halo mass $M$ normalized to the mass of the earth ($M_{\oplus}$) for different values of the free-streaming horizon, at a redshift $z=10$, and $T_{\rm RH}=10$ MeV. Clearly, for $M>M_{\rm RH}$, $\sigma(M)$ follows the rms density that would be found in a standard radiation-dominated universe with $\lambda_{\rm fsh}=0$. However, for $M_{\rm fsh}<M<M_{\rm RH}$, a power law growth $\sigma\left(M\right)\propto M^{-\left(n_{s}+3\right)/6}$ is observed, which becomes constant at scales below $M_{\rm fsh}$.
Figure 5: $\sigma(M)$ (left) and differential mass fraction (right) at $z=10$ are shown as a function of dark matter halo mass in units of earth mass($M_{\oplus}$), for different values of $\lambda_{\rm fsh}$. We set $T_{\rm RH}=10$ MeV which corresponds to $M_{\rm RH}=10.5\,M_{\oplus}$. The vertical dotted contours denote the value of $M_{\rm fsh}$ associated with the respective free-streaming horizons. The green dot-dashed curves represent the scenario with pure radiation domination and $\lambda_{\rm fsh}=0$.
@@
where $\delta_{c}\simeq 1.686$ is the critical value of linear overdensity for $z\geq 2$. In the right panel of Fig. 5 we show that the increase in the differential fraction significantly depends on the free-streaming horizon of the DM. In fact, we observe that if $\lambda_{\rm fsh}$ is below a threshold value $\lambda_{\rm fsh}^{c}$, the structures forming at scales smaller than $M_{\rm RH}$ in the presence of EMDE are guaranteed to show a boost in the differential fraction compared to the radiation dominated universe. The threshold $\lambda_{\rm fsh}^{c}\simeq 1.6$ pc primarily depends on $T_{\rm RH}$. Since we determine the threshold $\lambda_{\rm fsh}^{c}$ by comparing with a scenario where $\lambda_{\rm fsh}=0$ in the purely radiation-dominated universe, any $\lambda_{\rm fsh}<\lambda_{\rm fsh}^{c}$ in the case of EMDE ensures a boost in the structure formation.
Enhanced small-scale structures could amplify DM annihilation signals within sub-earth halos, providing a probe for a variety of DM models as well as non-standard cosmological epochs like EMDE. The dynamics of DM decoupling during EMDE determine the free-streaming horizon, which, when compared to the threshold horizon $\lambda_{\rm fsh}^{c}$, can help in identifying a viable range of microscopic model parameters compatible with enhanced structures, we defer further discussion to Sec. 4.
@@ 4.1 Model I: s-wave scattered scalar dark matter
The first model we consider has a scalar dark matter ($\phi_{\chi}$), interacting with scalar bath particles ($\phi_{\gamma}$) via a quartic vertex $(\lambda/4)\phi_{\chi}^{2}\phi_{\gamma}^{2}$. In general, $\phi_{\gamma}$ may represent either the Standard Model Higgs boson or any BSM scalar in the thermal bath. However, in the following, we assume that the mass of $\phi_{\gamma}$ is much smaller than all the relevant scales. This leads to an $s$-wave scattering $\phi_{\chi}\phi_{\gamma}\to\phi_{\chi}\phi_{\gamma}$, with momentum transfer rate
@@
where $m_{\chi}\in(20-1000)$ GeV is the typical mass for a WIMP DM.
@@
Figure 7: The allowed parameter space satisfying the present-day relic density of the DM is shown with distinct values of $T_{\rm RH}$ presented by coloured bands. The DM shows partial kinetic decoupling during the EMDE in the dark-shaded region, while in the light-shaded region, DM kinetically decouples after reheating. In the lime coloured region below the colour bands, DM produces non-thermally from the bath, while the light cyan region, where $T_{\rm RH}<10$ MeV, is excluded to avoid bounds from the BBN. The isocontours of the free-streaming horizon are shown by the white curves. Amplification of small-scale structures compared to the standard radiation-dominated cosmology occurs to the right of the cyan dot-dashed contour, which represents the cut-off value of the free-streaming horizon $\lambda^{c}_{\rm fsh}$.
@@
  • The DM was in thermal contact at the onset of the EMDE, and freezes-out prior to the end of reheating, the observed present-day relic density of the DM $\Omega^{0}_{\rm DM}h^{2}=0.12$ [2] is reproduced in the entire coloured region.
  • In the dark-coloured area, partial kinetic decoupling of the DM occurs, however, the sub-earth halo population amplifies below the cyan dot-dashed contour.
@@
A few comments regarding the impact of the temperature-dependent entropy injection during the EMDE are in order. First of all, the annihilation cross-section required to satisfy the relic density in this cosmological background is smaller than that required in a fully radiation-dominated universe, due to the dilution of the relic by the entropy injection. In fact, in the parameter space of Fig. 7 that shows freeze-out production of the DM, $\langle\sigma v\rangle_{\rm ann}$ ranges from $10^{-11}-10^{-15}$ GeV-2 depending on $m_{\chi}$, which is at least two orders of magnitude smaller than what would be required in the absence of EMDE. This reduction in the annihilation cross-section helps evade the strong bounds coming from the direct detection experiments. On the other hand, we emphasize that the required cross-section in this scenario is in principle larger than that in an EMDE with a constant rate of entropy injection ($\Gamma_{\phi}\sim{\rm const.}$). To give an idea, we estimate the ratio of $\langle\sigma v\rangle_{\rm ann}$, required to produce the same comoving number density of the DM after reheating in EMDEs with constant and temperature-dependent entropy injection rates as follows
@@
Here $x^{T}_{\rm fo}$ and $x^{c}_{\rm fo}$ denote the ratio $m_{\chi}/T_{\rm fo}$ in the two scenarios with $\Gamma_{\phi}\sim T$ and $\Gamma_{\phi}\sim{\rm const.}$, respectively. Assuming that the freeze-out occurs around $x_{\rm fo}\sim 20$ in both scenarios and neglecting the $T/m_{\chi}$ dependence in $\langle\sigma v\rangle_{\rm ann}$, we find that the annihilation cross-section required for the same amount of relic is approximately 50 times larger in the $\Gamma_{\phi}\sim T$ case compared to the $\Gamma_{\phi}\sim{\rm const.}$ case. However, we also point out that the present-day relic density of the DM, $\Omega^{0}_{\rm DM}h^{2}=0.12$ [2], cannot be satisfied by the quartic interaction during an EMDE with $\Gamma_{\phi}\sim{\rm const.}$ if $T_{\rm RH}\ll 10$ GeV [40]. Therefore, the temperature-dependent entropy injection accommodates a novel region of the parameter space where the relic density can be satisfied even with low reheating temperature $T_{\rm RH}\sim 100$ MeV and reduced annihilation cross-section compared to the radiation-dominated universe [^6] The constraints on $T_{\rm RH}$ derived in [95], based on fixed values for the WIMP mass, its annihilation cross-section, and the microhalo mass, cannot be directly applied to the present scenario due to the difference in the underlying assumptions, thus, we leave a detailed analysis of such constraints for future work..
The above discussion indicates the crucial impact of the background cosmological epoch on the production of DM. Moreover, in a large range of $m_{\chi}$, the DM with $s$-wave scattering shows partial kinetic decoupling during the EMDE. This is also unique for $\Gamma_{\phi}\sim T$ scenario, as $s$-wave partial amplitudes are not sufficient to kinetically decouple the DM in $\Gamma_{\phi}\sim{\rm const.}$ scenario, see Table 1. Thus, a broad region of the parameter space exhibits the correct freeze-out relic density with a low reheating temperature, leading to a reduced DM free-streaming horizon below the cut-off $\lambda^{c}_{\rm fsh}$. The resulting enhancement in small-scale structure formation at sub-earth mass scales offers new avenues to probe this scenario through DM annihilation signatures, provided the boost in the density of sub-halos can compensate for the reduced annihilation cross-section of the DM. However, alternate probes based on gravitational interaction of the DM such as pulsar timing arrays [103, 104, 95, 105], microlensing [106, 107], as well as the sub-halo DM interactions with cosmic rays [102], and sub-halo collisions with neutron stars could provide additional handles to study the early matter-dominated epoch.